OpenStudy (anonymous):

which equation represents the line that passes through the points (6 –3) and (–4 –9)

OpenStudy (mrnood):

OK - straight lines have an equation of general form y=mx+b where m is the gradient, and b is the y value when x=0 (known as the y intercept) Given 2 points you can work out the gradient in your case x1 = 6 y1 = -3 x2= -4 y2= -9 gradient is given by \[gradient = \frac{ y _{2}-y _{1} }{x _{2}-x _{1} }\]

OpenStudy (mrnood):

so now you have the value for the gradient m so the general formula is y = mx+b (and oyu know m) so the only unknown is b but oyu know 2 of the points which lie on the line - so pick one of them - lets say (6,-3) so y=-3, x=6 -3 = m(6) +b since you know m you can solve the above equation for b and then the answer is y=mx+b where you put in the values you calculated for m and b

OpenStudy (anonymous):

the answer choices are: y+4=3/5(x+9) y+4=5/3(x+9) y-3=3/5(x+6) y+3=3/5(x-6)

OpenStudy (mrnood):

I don't really care what the answer choices are. The information above is ALL you need to work out the answer for yourself. SO posting the answer choices AFTER I have given you all that information is a kind of indication that you don't really want to know HOW to do it after all!